OpenStudy (anonymous):

a bowl contains 40 balls numbered 1 through 40. if the numbers 1 through 10 are considered lucky and 2 balls are drawn from the bowl without replacement,find the probability that niether ball is lucky

OpenStudy (anonymous):

What are the odds that the first ball is not lucky?

OpenStudy (anonymous):

Probability of the first ball being unlucky = 30/40 = 3/4 Probability of the second ball being unlucky = 29/39 Probability of both = (Probability of first) * (Probability of the second)

OpenStudy (anonymous):

0.557692308